### v1

# AQORA: A Learned Adaptive Query Optimizer for Spark SQL

Abstract
Recent studies have identified two main approaches to improve query optimization: learned query optimization (LQO), which generates or selects better query plans before execution based on models trained in advance, and adaptive query processing (AQP), which adapts the query plan during execution based on statistical feedback collected at runtime. Although both approaches have shown promise, they also face critical limitations. LQO must commit to a fixed plan without access to actual cardinalities and typically rely on a single end-to-end feedback signal, making learning inefficient. On the other hand, AQP depends heavily on rule-based heuristics and lacks the ability to learn from experience. In this paper, we present AQORA, an adaptive query optimizer with a reinforcement learning architecture that combines the strengths of both LQO and AQP. AQORA addresses the above challenges through four core strategies: (1) realistic feature encoding, (2) query stage-level feedback and intervention, (3) automatic strategy adaptation, and (4) low-cost integration. Experiments show that AQORA reduces end-to-end execution time by up to 90% compared to other learned methods and by up to 70% compared to Spark SQL’s default configuration with adaptive query execution.
I Introduction
Query optimization plays a central role in database systems, determining the physical plan for a given query. Typically, this process takes place prior to query execution. The query optimizer selects a fixed execution plan based on estimated data statistics, analytical cost models, and a predefined set of transformation rules. Once execution begins, the query processing engine adheres strictly to the chosen plan, without the opportunity to revise it based on actual runtime behavior.
In response to the limitations of static, pre-execution query planning, subsequent research has evolved along two primary directions. One line of work focuses on developing learning-based algorithms that assist or replace the traditional query optimizer in making key planning decisions (such as join ordering, operator selection, or cardinality estimation) prior to execution beginning[1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. These learned query optimizers (LQOs) leverage historical execution data to enhance decision-making and generalize to unseen queries. In parallel, another line of research shifts part of the optimization responsibility into the execution phase, enabling dynamic plan adaptation based on actual runtime statistics[13, 14, 15, 16, 17, 18, 19, 20]. This paradigm, known as adaptive query processing (AQP), has been increasingly adopted by modern engines such as Spark SQL[21, 22], allowing certain decisions to be deferred until runtime.
While both approaches have significantly advanced the state of query optimization, they each face inherent limitations stemming from their core methodology. For LQOs, their optimization is confined to the pre-execution phase, which gives rise to two fundamental challenges:
C1: Lack of runtime information. Pre-execution query optimization requires the system to make planning decisions without access to essential runtime information, such as the actual cardinalities of intermediate results. Although researchers have proposed various encoding methods and predictive models to improve pre-execution estimation, these approaches still often fail to consistently outperform traditional rule-based optimizers such as PostgreSQL’s planner [23].
C2: Lack of fine-grained feedback. Most LQOs receive only a single final feedback metric (e.g., end-to-end execution time) for each query, even though they make multiple interdependent optimization decisions before execution. This coarse-grained signal fails to identify which specific decision led to performance changes. More critically, LQOs cannot observe intermediate execution states and revise their decisions based on actual runtime behavior.
On the other hand, prior work [24] has shown that even simple AQP techniques can, in some cases, match or even outperform more sophisticated LQOs. However, AQP methods also face two fundamental challenges:
C3: Lack of flexible adaptation. Most existing AQP methods rely on static, rule-based heuristics to trigger runtime adjustments, such as switching join strategies or enabling dynamic repartitioning. These rules are often tuned manually via parameters set by DBAs, which limits their flexibility and generalization to unseen workloads. Moreover, such methods rarely consider join reordering, despite it being one of the most impactful aspects of query optimization.
C4: Lack of practical deployability. Although AQP methods offer many valuable ideas, only a small number have been implemented in real-world database systems. A main barrier is the inherent complexity of combining planning and execution, which complicates testing, debugging, and performance tuning. Introducing learning-based mechanisms in this context further increases system complexity, making reliable deployment even more challenging.
This leads to a fundamental question: can we design a system that effectively combines the strengths of pre-execution learning with in-execution adaptation? Fig. 1 illustrates the design space of query optimizers along two key axes: when optimization occurs (pre-execution vs. in-execution) and how decisions are made (rule-based vs. learning-based). Most LQOs, such as AutoSteer [9] and Lero [10], perform the optimization in the pre-execution phase and make decisions without access to runtime feedback. In contrast, AQP methods, such as Spark SQL’s Adaptive Query Execution and runtime filters, perform limited rule-based adjustments during execution. As shown in the figure, our proposed system, AQORA (Adaptive Query Optimizer with Reinforcement learning Architecture), is the first to reside in the top-right quadrant: a learned query optimizer that performs in-execution plan adaption on Spark SQL. It integrates with Spark’s runtime engine via a lightweight planner extension and uses an actor-critic reinforcement learning model to make informed decisions during execution, leveraging query stage-level feedback for fine-grained supervision. It is designed as a plug-and-play module, requiring minimal changes to the underlying system, and can be easily integrated into Spark SQL.
Figure 1: AQORA is the first learned query optimizer achieving execution-time query plan optimization on Spark SQL.
AQORA addresses the aforementioned challenges through four key strategies:
S1: Runtime plan feature representation. AQORA removes the constraint of making all optimization decisions before execution begins. Instead, it dynamically collects runtime signals, such as actual cardinalities and execution statistics, and uses them as query plan features. This enables the model to make more informed and adaptive decisions based on real-time execution context. Compared to approaches that rely solely on pre-execution estimation [25, 26], this design also reduces the reliance on handcrafted features and complex estimators while improving robustness and accuracy.
[…]
S3: Learning-based adaptive strategies formulation. Instead of relying on fixed, rule-based heuristics, AQORA’s decision model automatically learns its optimization strategies from feedback collected during query execution. This data-driven approach enhances generalization across diverse query workloads without requiring manual tuning. To enhance learning stability and final performance, AQORA adopts a curriculum learning strategy: it starts with simpler decisions and gradually increases the complexity, enabling the model to learn more effective strategies in a step-by-step manner. In addition, AQORA supports a two-phase optimization mechanism, allowing strategies learned during execution to be reused in the planning phase of future queries.
S4: Lightweight and seamless system integration. Unlike prior approaches [10, 27] that require intrusive source code modifications, AQORA adopts a non-intrusive, plugin-style architecture. It can be enabled through simple configuration adjustments without altering the underlying query engine’s codebase. This plug-and-play design significantly reduces integration overhead and enables seamless deployment through minimal configuration.
Our contributions are summarized as follows:
  • We propose AQORA, an adaptive query optimizer for Spark SQL that integrates reinforcement learning into Spark’s adaptive query execution framework. AQORA comprises a decision model based on actor-critic learning and a planner extension that enables online plan modification during execution.
  • We adopt a lightweight decision model built upon the TreeCNN architecture [28], which learns optimization actions directly from stage-level feedback collected during query execution. By using true cardinalities as input features (and avoiding sample bitmaps or histogram encodings), the model achieves better query performance with a simplified architecture.
  • We design a plug-and-play planner extension built upon the extensibility interfaces of modern query engines, enabling non-intrusive integration with existing deployments. The extension intercepts runtime execution plans, interacts with the decision model, and applies optimization actions such as join reordering.
  • We conduct extensive experiments on three public benchmarks. Results show that AQORA reduces end-to-end execution time by up to 90% compared to other learned methods and by up to 70% compared to Spark SQL’s default configuration with adaptive query execution.
II Related Work
In this section, we review related work, including LQOs and previous AQP methods. Since the pioneering work of LEO [29] in 2001, which introduced the concept of a learning-based query optimizer, recent studies on LQO have evolved into two main directions: plan-constructors and plan-steerers.
Plan-constructor
Plan-constructor methods aim to build query plans directly from scratch. These approaches enable the discovery of novel and potentially more efficient plans that lie beyond the capabilities of traditional optimizers. Neo [1] pioneers this category by leveraging past queries to progressively construct improved query plans, using the native optimizer as a weak oracle. Balsa [2] further improves the convergence efficiency of plan learning by employing enhanced initialization and training techniques. LOGER [3] enhances interpretability in learned planning by detecting detrimental operators and adjusting reward weighting accordingly. GLO [4] improves generalization to unseen workloads by clustering database statistics and learning generalized policies across these clusters. However, constructing plans from scratch suffers from two fundamental limitations: 1) the initial search often begins with poorly performing plans; and 2) the search space is extremely large, making efficient convergence challenging. In contrast, AQORA addresses these issues by starting from a reasonable initial plan generated by the native optimizer and learning to make incremental adjustments within a constrained action space.
Plan-steerer
Plan-steerer methods improve query plans by exploiting the internal mechanisms of the native optimizer. Instead of constructing plans from scratch, they guide the optimizer to explore multiple candidate plans and select the one with the lowest estimated cost. For example, Bao [5], DeepO [8], SCOPE [6], FASTgres [7], and AutoSteer [9] generate different query plans by toggling rule-based heuristics (i.e., hint sets) within traditional optimizers. HybridQO [11] explores alternative join orders by injecting leading hints. Lero[10] modifies cardinality estimates in the native optimizer to produce diverse plans. LEON [12] enumerates sub-plans directly using the native optimizer, while FOSS [30] refines plans after initial optimization by fixing join orders and methods. However, these methods rely heavily on the internal mechanisms of specific DBMS-native optimizers. As a result, the effectiveness of some of these methods is limited when applied to different DBMSes.
Adaptive query processing
AQP has been extensively studied to enable query plan adjustments during execution based on runtime feedback [13, 14, 15, 16, 18, 17, 19, 20, 27]. For example, Lookahead Information Passing [19] generalizes semi-join techniques to enhance the pipelining of equijoins. Adaptive Join Algorithm [31] dynamically switches between hash join and nested loop join during runtime. UDAO [27] adjusts parameters at multiple levels of granularity throughout query execution. RPT [32] extends predicate transfer with efficient semi-join reduction and join order validation to achieve provably robust join ordering for acyclic queries. In contrast to these methods, AQORA introduces the first learning-based AQP framework capable of continuously improving from experience and adaptively optimizing joins at runtime.
III Preliminary
In this section, we formulate the problem of query optimization and provide background on Spark SQL’s current approach.
III-A The Problem of Query Optimization
Query optimization involves many considerations, including predicate pushdown, expression simplification, and join optimization, etc. In this paper we focus on two core decisions: (1) determining the optimal join order and (2) selecting appropriate physical operators.
Consider a query $Q$ that joins $n$ tables $t_{1},t_{2},\cdots,t_{n}$}. If we restrict plan space to left-deep trees (Fig. 2(a)), there are up to $n!$ possible join orders. Allowing bushy trees (Fig. 2b) enlarges the search space to (2n2)!/(n1)!(2n-2)!/(n-1)! [33].
(a) left-deep join plan
(b) bushy join plan
Figure 2: In a left-deep tree plan, the left child of any join operation is a subtree, and the right child must be a leaf node. In contrast, bushy tree plans allow both children of any join operation to be subtrees
Join order and operator selection are critical factors in determining query execution efficiency. The core challenge of query optimization is to find a query plan $p$ that specifies both a join order and a set of physical operators, while minimizing the overall execution cost, which can be expressed as follows:
$\arg\min_{p}C(p)=C_{\text{plan}}(p)+C_{\text{execute}}(p)$
Here, $C_{\text{plan}}$ denotes the time spent generating candidate plans and identifying the final plan $p$, while $C_{\text{execute}}$ represents the actual execution time of plan $p$.
III-B Cost Based Optimization
Spark SQL’s cost-based optimizer (CBO) selects join orders using a dynamic programming algorithm that relies on cardinality estimates. This approach performs reasonably well for simple queries, where even with inaccurate cardinality estimates, suboptimal join orders still keep the planning cost ($C_{\text{plan}}$) within acceptable bounds. However, it encounters significant challenges when handling complex queries with many joins. The exponential time complexity of dynamic programming leads to rapid growth in optimization time, often resulting in both prolonged $C_{\text{plan}}$ and suboptimal plan selection. As shown in Fig. 3, $C_{\text{plan}}$ can become the dominant component of total cost as the number of joins increases. In fact, for JOB query 29a, the execution time shows little difference between enabling and disabling CBO, whereas the optimization time dominates when CBO is enabled. While we acknowledge that this issue may be specific to Spark SQL’s CBO implementation, our goal is not to replace or circumvent CBO but to provide a practical mechanism that dynamically mitigates such performance pitfalls.
Figure 3: Execution and optimization times of selected JOB queries with and without CBO.
CBO determines which join operator to use during physical optimization based on cardinality estimates. A key configuration is autoBroadcastJoinThreshold (BJT), which controls when broadcast joins are applied. If either side of the join is estimated to be smaller than BJT, Spark uses a broadcast hash join; otherwise, it falls back to shuffle joins such as sorted merge join.
However, this approach can lead to two problems. First, inaccurate cardinality estimates may result in selecting an inappropriate join type. Second, if the actual data size exceeds BJT but the estimate is incorrect, Spark may attempt to broadcast a large table, potentially causing worker crashes due to memory overload. To prevent such crashes, administrators often set BJT to a conservative value. While this safety measure enhances stability, it also limits the performance benefits offered by broadcast joins.
III-C Adaptive Query Execution
Spark SQL 3.0 (2020) introduced Adaptive Query Execution (AQE), which is an implementation of AQP and aims to tackle the challenges of cardinality estimation. As Xue et al. [22] claimed, maintaining accurate statistics in distributed systems like Spark is more difficult than in traditional databases because the continuous flow of data makes real-time statistics maintenance impractical. However, abandoning cardinality estimation is not feasible, as optimization without statistical guidance tends to perform even worse. AQE’s solution is to defer certain decisions (like determine join operators): it starts with rough estimates and then re-optimizes the query during execution using actual runtime metrics such as shuffle partition sizes, as illustrated in Fig. 4.
Figure 4: This query performs three joins: after the first join $t1\bowtie t2$ produces only 5 rows, AQE switches the second join from SMJ to BHJ; after the second join $(t1\bowtie t2)\bowtie t3$ produces 100,000 rows, AQE switches the third join from BHJ to SMJ.
AQE leverages runtime statistics to improve broadcast join decisions and adjust physical operators. It also performs specific optimizations such as merging shuffle partitions and fixing skewed joins. However, AQE cannot modify the initial join order determined during logical optimization and lacks the capability to learn from past executions.
IV System Overview
In this section, we introduce the main components of AQORA and its basic workflow.
As shown in Fig. 5, AQORA consists of two core components: a planner extension and a decision model. The workflow proceeds as follows:
Figure 5: An overview of AQORA, which consists of two main components: (1) a decision model that generates optimization actions, and (2) an AQE planner extension that applies these actions and provides feedback to the decision model.
  1. 1.
    When Spark SQL’s AQE is triggered during execution, the planner extension extracts the current query plan along with actual cardinality statistics and forwards it to the decision model. This process can also be triggered during the query optimization phase, in which case actual cardinalities are not available.
  2. 2.
    The decision model analyzes the input plan and returns optimization actions (e.g., swapping the join order of leaf nodes).
  3. 3.
    The planner applies these actions to the ongoing query plan and resumes execution using the optimized version.
AQORA iteratively performs steps 1)-3) until the query completes and then performs step 4). We formalize this $k$-step process as a trajectory:
$\tau=(s_{0},a_{0},r_{1},s_{1},\cdots,a_{k-1},r_{k},s_{k})$
where each state $s_{i}$ consists of a partially executed plan $p_{i}$ and its corresponding cardinalities $c_{i}$, while $a_{i}$ denotes the action taken at step $i$, and $r_{i+1}$ represents the shaping reward obtained after taking action $a_{i}$.
4) After the query completes, the decision model replays the collected trajectory to learn from past decisions and improve its optimization strategy for future queries.
V Decision Model
In this section, we describe the decision model in detail.
V-A The Actor-Critic Framework
We adopt the actor–critic framework to fully exploit the frequent feedback signals from AQE, which promotes faster convergence and allows us to observe the critic’s reward or penalty for each action taken by the actor.
V-A1 State, Action, and Return
To introduce the framework, we formally define the three main components of the Markov decision process for query optimization:
State
Each state $s_{t}=(p_{t},c_{t})$ consists of:
  • $p_{t}$: the partial executed plan at time step $t$;
  • $c_{t}$: the true cardinalities of completed plan segments, obtained from AQE statistics.
Action Space
AQORA supports four types of actions: turning on/off CBO, manipulating join orders, setting broadcast join hints on leaf nodes to enforce broadcast joins, and performing a no-op. The notations and definitions of these actions are summarized in Tab. I:
TABLE I: Action Space of AQORA’s Decision Model
Notation Definition
cbo(1/0)(1/0) turn on/off CBO
swap(i,j)(i,j) swap the $i$-th and $j$-th leaf node
[…]
broadcast(i)(i) broadcast the $i$-th leaf node
no-op do nothing
The action space of AQORA is carefully designed to balance flexibility and tractability in adaptive query optimization. The first type (CBO action) is included because the quality of the initial query plan significantly impacts the subsequent optimization trajectory. Dynamically controlling this setting allows the agent to make better trade-offs between optimization quality and performance cost.
Once an initial plan is chosen, AQORA can refine it through two kinds of structural modifications: adjusting the join order and modifying the join method. The lead action provides a more practical and efficient mechanism for guiding join order, as most beneficial changes can be achieved by repositioning a single leaf node to the front.
For join method selection, the key issue is the overly conservative application of BJT. To address this, AQORA provides the broadcast action, which forces a broadcast join on a selected leaf node. While this strategy introduces some risk, it enables effective correction of conservative join method choices.
Finally, the no-op action allows the model to intentionally defer decision and wait for more runtime information.
Return
Based on empirical validation, we define the return as:
$R(\tau)=\sum_{i=1}^{k}\gamma^{i-1}r_{i}-\sqrt{T_{\text{execute}}(\tau)}$
where $T_{\text{execute}}$ represents the total execution time of $\tau$ (in seconds). The square root operation helps mitigate reward scale variations across different queries. If query execution fails, $R(\tau)$ is assigned a substantial negative penalty (e.g., $\sqrt{300}$). The intermediate reward $r_{i}$ is defined as the negative value of the additional number of shuffles introduced after taking action $a_{i}$, normalized by 10. This definition naturally provides positive feedback when the model selects no-op (no additional shuffle is introduced), encouraging safe behavior during early training, while penalizing actions that introduce extra shuffles and discouraging costly changes late in the optimization process. The discount rate $\gamma(<1)$ ensures that the accumulated intermediate rewards $r_{i}$, even over many steps, do not dominate the overall cost $\sqrt{T_{\text{execute}}}$.
V-A2 Policy and State Transition
The actor network implements a stochastic policy $\pi_{\theta}(a|s)$, which maps states to action probabilities through softmax normalization:
$\pi_{\theta}(a|s)=\begin{cases}P_{1},\quad&\text{if}\ a=a_{1}\ (\text{e.g.,swap}\ (1,2))\\ P_{2},\quad&\text{if}\ a=a_{2}\ (\text{e.g.,broadcast}\ (1))\\ &\vdots\\ P_{m},\quad&\text{if}\ a=a_{m}\ (\text{e.g., no-op})\end{cases}$
where $\sum_{i=1}^{m}P_{i}=1$. The probabilities $P_{i}$ are computed by the actor network’s parameters $\theta$ via softmax normalization over its output logits. This probabilistic approach enables exploring alternative optimization paths while maintaining likely good choices.
After the actor takes an action $a$, the current state $s$ transitions to the next state $s^{\prime}$. The state transition probability $P(s^{\prime}|s,a)$ reflects the uncertainty in adaptive execution:
[…]
The uncertainty arises from AQE’s non-deterministic triggering: the execution engine may complete a varying number of stages between optimizations. As a result, applying the same action $a$ in state $s$ can lead to different successor states $s^{\prime}$.
V-A3 State Value and Action Value
The state value $v_{\pi}(s)$ quantifies the expected long-term utility of being in state $s$. The critic evaluates this value based on the current partial plan and runtime statistics. For example, if a join operation involves high cardinalities on both sides—often indicating expensive data shuffling—the critic may assign a lower $v_{\pi}(s)$ compared to states with smaller intermediate results. Note that this pattern isn’t universal—while some large joins may be unavoidable for optimal performance, it serves as an intuitive starting point. Through training, the critic learns to refine these initial heuristics into precise value estimates.
Also notice that the state value $v_{\pi}(s)$ depends not only on the current plan $s$ but also on the actor’s policy $\pi$. Even when reaching a promising state $s$, a suboptimal policy may select actions that degrade subsequent states. For example, a policy favoring broadcast joins for large tables could turn $s$ (with moderate cardinalities) into $s^{\prime}$ (with memory pressure), lowering $v_{\pi}(s)$ despite $s^{\prime}$’s inherent potential. The critic’s estimated value $v_{\phi}(s)$ approximates $v_{\pi}(s)$, learning to anticipate these policy-induced outcomes through exposure to the actor’s decision patterns.
The action value $q(s_{t},a_{t})$ measures the impact of action $a_{t}$ on optimization potential. Let $v_{\phi}(s)$ denote the critic’s value estimate with parameters $\phi$. We compute:
$q(s_{t},a_{t})=r_{t+1}+v_{\phi}(s_{t+1})-v_{\phi}(s_{t})$
where $s_{t+1}$ is the state after performing $a_{t}$. For instance, swapping larger tables early (action $a_{t}$) might decrease $v_{\phi}(s_{t+1})$ due to increased shuffle costs, yielding negative $q(s_{t},a_{t})$. Conversely, broadcasting a small table (action $a_{t}^{\prime}$) could improve $v_{\phi}(s_{t+1})$, producing positive action value. This differential signal guides the actor’s policy updates.
V-A4 Training Algorithm
For the actor, our goal is to maximize the expectation of $q(s_{t},a_{t})$ for each step. So we use the clipped PPO loss [34] with an entropy factor to encourage explorative actions:
$L^{\text{clip}}=-\frac{1}{k}\sum_{t=0}^{k-1}\left[\min\left(\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})},\text{clip}\left(\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}(a_{t}|s_{t})}},1\pm\epsilon\right)\right)q_{t}\right]$
$L^{\text{entropy}}=\frac{1}{k}\sum_{t=0}^{k-1}\pi_{\theta}(a_{t}|s_{t})\log\pi_{\theta}(a_{t}|s_{t})$
$L^{\text{actor}}=L^{\text{clip}}+\eta L^{\text{entropy}}$
For the critic, our goal is to minimize the bias of the evaluated state value $v_{\phi}(s)$ and the ground truth state value $v_{\pi}(s)$, so we use the MSE loss function to optimize the critic’s parameters $w$:
$L^{\text{critic}}=\frac{1}{k}\sum_{i=0}^{k-1}(v_{\phi}(s_{t})-v_{\pi}(s_{t}))^{2}$
We observe that the policy gradient for updating the actor depends on $q$ (calculated from the critic’s state values), while the critic’s update gradient relies on $v_{\pi}$ (determined by the actor’s current policy).
As shown in Alg. 1, this bootstrapping interaction allows both components to progressively improve through repeated optimization iterations, where $\gamma$ is set to 1 for simplicity.
Algorithm 1 The training algorithm in AQORA
1:Initial policy parameters $\theta$, initial value parameters $\phi$
2:Clip parameter $\epsilon$, training epochs $e$
3:Collect trajectory $(s_{0},a_{0},r_{1},s_{1},\cdots,a_{k-1},r_{k},s_{k})$ using current policy $\pi_{\theta}$
4:Get actual state values $V_{\pi}:(v_{\pi}(s_{0}),v_{\pi}(s_{1}),\cdots v_{\pi}(s_{k}))$, $v_{\pi}(s_{i})=\sum_{j=i+1}^{k}r_{j}+\sqrt{T_{\text{execute}}}$
5:Get inferred state values $V_{\phi}:(v_{\phi}(s_{0}),v_{\phi}(s_{1}),\cdots v_{\phi}(s_{k}))$, $v_{\phi}(s_{i})$ is inferred from critic net’s initial parameters $\phi$
6:Compute action values $Q=(r_{1}+v_{\phi}(s_{1})-v_{\phi}(s_{0}),r_{2}+v_{\phi}(s_{2})-v_{\phi}(s_{1}),\cdots,r_{k}+v_{\phi}(s_{k})-v_{\phi}(s_{k-1}),0)$
7:Record old action probabilities $\pi_{\text{old}}=\pi_{\theta}(a_{t}|s_{t})$ inferred from actor net’s initial parameters $\theta$
8:for epoch =1=1 to $e$ do
9:  Compute new probabilities $\pi_{\theta}(a_{t}|s_{t})$
10:  Calculate probability ratio $r_{t}=\pi_{\theta}(a_{t}|s_{t})/\pi_{\theta_{\text{old}}}(a_{t}|s_{t})$
11:  Compute actor loss: $L^{\text{actor}}$
12:  Compute critic loss: $L^{\text{critic}}$
13:  Update $\theta\leftarrow\theta-\alpha\nabla_{\theta}L^{\text{actor}}$
14:  Update $w\leftarrow w-\beta\nabla_{w}L^{\text{critic}}$
15:end for
From the above analysis, we see that the critic evaluates each individual action taken by the actor, providing feedback that is $k$ times denser than that of end-to-end reward methods. When the actor makes a suboptimal decision (e.g., selecting an inefficient join order), it receives immediate feedback via the critic’s updated state value, without needing to wait for the entire query to complete. Which means even if subsequent actions compensate for earlier mistakes (e.g., correcting a bad join order later in the plan), the initial suboptimal action still receives its own critical feedback. This ensures that every action contributes to policy updates, not just the final outcome.
V-B Model Design
This section presents the implementation details of the actor’s policy $\pi_{\theta}(s)$ and the critic’s value function $v_{\phi}(s)$.
When the decision model receives a partial plan $s$, it first compresses the plan tree and encodes each of its nodes into embedding vectors. The resulting vectorized tree is then processed separately by the actor and critic networks. The actor generates a probability distribution $\pi_{\theta}(s)$ over the set of valid actions, from which an action is selected using masked probabilities to filter out illegal operations. Meanwhile, the critic produces a scalar estimate $v_{\phi}(s)$, representing the state’s optimization potential.
V-B1 Tree Compressing
While nodes such as sorts and aggregates do affect query performance, we exclude them from the model’s input features to focus solely on join operations (Fig. 6(1)). This approach, though risky for other LQOs relying on full plan details, becomes viable because AQORA’s access to real-time cardinalities compensates for the lost information.
Figure 6: Tree compressing and feature encoding.
V-B2 Feature Encoding
The compressed plan tree is transformed into a vector tree, as shown in Fig. 6(2). Each node $u$ is encoded as a vector by concatenating three components:
[…]
with the following encoding rules:
  • Each node is assigned a type, such as a join or a broadcast query stage. AQORA encodes the node type as a one-hot vector type(u)(u). In a simplified example shown in Figure 6, if $u$ is a leaf node, then type(u)=(0,1)(u)=(0,1); if $u$ is a join node, then type(u)=(1,0)(u)=(1,0)
  • $\text{table}(u)$ is a binary (0/1) vector indicating which tables participate in node $u$, as shown in Fig. 6. It encodes the join sequence leading to node $u$, and during AQE, even leaf nodes may touch multiple tables.
[…]
This encoding strategy ensures that AQORA remains operational under data updates and schema changes. Specifically, the operator type $\text{type}(u)$ is fixed, while the intermediate cardinality $\text{card}(u)$ is collected online during execution rather than relying on precomputed statistics. Even when new tables are introduced, the encoding remains valid, with the corresponding positions taking a default value of 0.
The encoded vector tree is processed by AQORA’s actor and critic networks.
We choose TreeCNN for its low optimization overhead and superior performance after being trained on a sufficient number of queries (2400 on ExtJOB).
V-B3 Masked Output
To prevent invalid actions such as swapping nonexistent nodes, we apply masking to the actor’s output. For a workload with at most $n$ tables, the dimension of the actor’s output is given by:
$d_{\pi_{\theta}(s)}=2+(n-1)+C_{n}^{2}+n+1$
for switching on/off CBO, n1n-1 possible leads (if we only consider lead a single table each time), $C_{n}^{2}$ possible swaps (each concrete action instance is treated as a distinct action, e.g., $\text{swap}(4,7)$), $n$ possible tables to broadcast, and 1 no-op. For example, when processing a partially executed plan with 3 leaf nodes, we create a binary mask as follows:
$\begin{split}M=(&\underbrace{0,0}_{\text{cbo(1), cbo(0), }},\underbrace{1,1,0,\cdots,1,0}_{\text{swap(1,2), swap(1,3), swap(2,3)}},\\ &\underbrace{1,1,0,\cdots}_{\text{lead(2), lead(3)}},\underbrace{1,1,1,0,\cdots}_{\text{broadcast(1), broadcast(2), broadcast(3)}},\underbrace{1}_{\text{no-op}})\end{split}$
However, to achieve smoother policy convergence, we implemented a curriculum learning framework that requires masking out additional actions at different training stages. Initially, the model is only allowed to choose whether to enable or disable CBO, focusing on a simple binary decision task. During the mid-training phase, we progressively lift the masking on runtime plan adjustment actions; as these decisions are based on actual runtime statistics, they are generally easier for the model to learn. In the final training stage, all restrictions are removed except for masking invalid actions, allowing the model to explore the full action space and refine its decision-making capabilities.
The final action probabilities are calculated by normalizing the masked output:
$\pi_{\text{final}}(s)=M\odot\pi_{\theta}(s)$
VI Planner Extension
Our planner extension operates through two core mechanisms: capturing and sending the current partial plan to the decision model and applying optimization actions returned by the actor to modify the query plan. As detailed at the end of Sec. VI-B1, these mechanisms support dynamic exploration of bushy joins through iterative plan modifications at runtime.
VI-A Send the Extracted Plan
The planner extension introduces a new optimization rule into existing rule set. This rule is triggered after all default (adaptive) optimization rules have been applied. Upon activation, it first checks whether the maximum number of allowed optimization steps has been reached. If not, it extracts the current partial execution plan along with available runtime cardinalities from each completed stage. The extracted plan and cardinalities are then sent to the decision model, after which the system awaits optimization instructions.
The setting of the maximum number of optimization steps plays a critical role in overall performance: higher values allow more optimization attempts and richer feedback signals, but they also require consistently accurate decisions across diverse scenarios. By default, we empirically set the maximum number of optimization steps (both before and during execution) to 3.
VI-B Inject Plan
The optimization instructions may involve five types of operations, as shown in Tab. I. While actions such as enabling CBO and no-op operations are straightforward to handle, different strategies are required for processing join reorder and broadcast instructions.
VI-B1 Plan Transformation
Swapping instructions are denoted as swap(i,j)(i,j), indicating the exchange of positions between the $i$-th and $j$-th leaf nodes. However, if the opposite sides of the join of these two leaf nodes lack compatible join keys, executing the swap would result in a Cartesian product. To mitigate this, we design Alg. 2 to perform controlled node swapping. Leading actions follow a similar strategy.
Algorithm 2 Swapping Algorithm
1:Plan $p$ with leaf nodes $l_{1},l_{2},\cdots,l_{n}]$ and join conditions $\{c_{1},c_{2},\dots,c_{m}\}$, swap instruction swap(i,j)(i,j) or lead instruction lead(i)(i)
2:New plan $p^{\prime}$ if the $i$-th and the $j$-th leaf nodes can be swapped or the $i$-th leaf node can be leaded, otherwise original plan $p$
[…]
5:Initialize $p^{\prime}\leftarrow l_{1}$
6:for $k=2$ to $n$ do
7:  Find condition $c\in C$ that cannot be evaluated separately on $p^{\prime}$ or $L^{\prime}_{k}$ but can be evaluated on $p^{\prime}\bowtie L^{\prime}_{k}$
8:  if $c$ exists then
9:   $p^{\prime}\leftarrow$ buildJoin($p^{\prime},L^{\prime}_{k},c$)
10:  else
11:   return $p$
12:  end if
13:end for
14:return $p^{\prime}$
We observe that applying swap (or lead) operations during a single execution can gradually transform the join plan into a bushy structure. As an illustrative example, consider the swap operation:
$((t_{1}\bowtie t_{2})\bowtie t_{3})\bowtie t_{4}\xrightarrow{\text{swap}((t_{1}\bowtie t_{2}),t_{4})}(t_{1}\bowtie t_{2})\bowtie(t_{3}\bowtie t_{4})$
This example highlights AQORA’s capability to explore bushy join plans through execution-time transformations. By applying the swap operation between the completed $t_{1}\bowtie t_{2}$ subtree and $t_{4}$, the system incrementally constructs a bushy plan during execution.
VI-B2 Join Operator Selection
Upon receiving a broadcast instruction, the planner extension performs a bottom-up traversal of the query plan. It first locates the corresponding join operator within the execution plan hierarchy, then annotates the identified join operator with a BROADCAST hint on the appropriate (left/right) side. This modification instructs Spark SQL to prioritize the broadcast join strategy for the specified join operation.
VII Experiments
In this section, we conduct extensive experiments to validate the effectiveness of AQORA.
VII-A Experiment Setup
VII-A1 Environment
All experiments were conducted on Spark 3.5.4 with 6 executors, each configured with 6 executor cores and 20 GB of executor memory. The decision model was trained and executed for inference on an NVIDIA A40 GPU.
VII-A2 Benchmarks
We selected three standard benchmarks, following prior work [4, 11].
JOB
[35] employs the real-world IMDb dataset with 21 tables related to movies and actors. It includes 113 complex SQL queries generated from 33 templates, each involving joins across 4 to 17 tables.
Extended JOB
[1] (referred to as ExtJOB) is an extension of the JOB benchmark, consisting of 24 queries derived from 12 distinct templates. Compared to the original JOB queries, ExtJOB features entirely different join graphs and predicates, making it a more challenging benchmark[2].
STACK
[5] comprises 10 tables from Stack Exchange datasets. Its 16 query templates generate analytical queries joining 4-12 relations. We excluded templates #4,7 (with arithmetic operations in join conditions) and #9,10 (containing unsupported predicate subqueries)—limitations shared by most existing LQOs. We plan to address these restrictions in the future.
VII-A3 Baselines
We compare AQORA against three state-of-the-art baselines:
Spark SQL
Combined with runtime filters and dynamic join selection, Spark SQL’s default configuration with AQE represents a strong baseline for adaptive query optimization, and it is also the most common industrial practice.
Lero
Lero [10] employs a learned comparator to evaluate join plans generated by traditional optimizers under different cardinality estimations, then chooses the best one among them.
AutoSteer
AutoSteer [9] systematically evaluates all available optimization rules in Spark SQL by disabling them to assess their impact on the current plan. It then constructs a collection of rules to disable for performance gains using greedy search.
We adopt the implementations of Lero and AutoSteer from their authors[36, 37]. We exclude other notable optimizers like Bao [5] and Balsa [2] primarily due to their incompatibility with Spark SQL, and we exclude UDAO[27] because it operates in a disjoint optimization space of parameter tuning.
VII-A4 Settings
Data Generation
For JOB and ExtJOB, we expanded the original IMDb dataset to ten times its original size. For STACK, we utilized the benchmark’s original archived dataset. All datasets were converted to Parquet format. Note that this experimental setup already imposes substantial memory pressure on our testing environment, as an bad query plan can easily lead to out-of-memory errors and execution failure.
Query Generation.
We generated 1,000 training queries using each benchmark’s templates. For each template, randomized predicate conditions were introduced while preserving the original join structure.
For JOB and ExtJOB, we exclusively used their original 113 and 24 queries as their test set, respectively. For STACK, the test set contains 10 additional queries per template generated in the same manner.
Evaluation Approach.
We conduct both static and dynamic evaluations. For the static evaluation, each optimizer is trained on the training set of a given benchmark until convergence, and then its performance is assessed by executing the corresponding test queries. For the dynamic evaluation, we train the optimizer on one dataset and then test it either on the test set of another dataset or on the same dataset after dynamically modifying the underlying data through updates.
Key Metrics
We evaluate performance using three core metrics: (1)End-to-end execution time $C$: total time required to run all test queries; (2)Optimization time $C_{\text{plan}}$: time spent generating optimization decisions; and (3)Raw execution time $C_{\text{execute}}$: time spent on purely executing the query plan. Each query’s execution time is capped at 300 seconds; queries that exceed this limit or fail due to out-of-memory are recorded as 300 seconds.
VII-B Query Performance
Fig. 7 shows the end-to-end execution time, optimization time, and raw execution time for all test queries across different methods.
Figure 7: Query performance on three benchmarks, in seconds. Black numbers outside bars represent end-to-end query execution time, while numbers in dark and light bars represent raw query time and optimization cost respectively.
VII-B1 End-to-End Execution Time
AQORA outperforms all three baselines across the three benchmarks. On the JOB benchmark, AQORA reduces end-to-end execution time by 47.3%, 90.1%, and 60.6% compared with Spark SQL’s default configuration, AutoSteer, and Lero, respectively. For ExtJOB, AQORA achieves reductions of 20.0%, 26.2%, and 38.5% against the same baselines. On STACK, AQORA reduces execution time by 73.3%, 17.2%, and 73.1% compared to Spark SQL’s default, AutoSteer, and Lero, respectively.
These results highlight the advantage of AQORA’s novel idea of optimizing queries during execution, surpassing traditional optimizers and other state-of-art methods.
VII-B2 Optimization Time
Spark SQL’s default optimization strategy incurs no additional overhead, as it directly executes the join order specified in the input SQL text. Consequently, we exclude its optimization time from our comparisons.
For learned methods such as Lero and AutoSteer, the optimization cost primarily stems from generating candidate plans. Lero achieves this by adjusting cardinality estimates at various levels of the query plan, producing over 10 candidate plans per query on average, with each EXPLAIN command taking about 10.1 seconds. AutoSteer, in contrast, generates more than 2 plans per query by enabling or disabling different optimization rules, and each EXPLAIN command takes around 3.3 seconds. This substantial overhead can, in some cases, lead to worse overall performance compared to Spark SQL’s default configuration, even when better query plans are discovered.
AQORA incurs significantly lower optimization costs, as it avoids the need to execute multiple EXPLAIN queries to enumerate candidate plans. Most of its overhead arises from model inference and the execution of Alg. 2, both of which are lightweight. However, AQORA introduces a new type of optimization overhead in distributed engines: modifying partially executed plans at runtime may trigger additional shuffle operations. Despite this, our experiments show the average optimization cost remains below 320 ms per query.
VII-B3 Raw Execution Time
The results show that the raw execution time of AQORA is comparable to that of Lero across all benchmarks. This indicates that modifying query plans during execution can be as effective as generating and evaluating multiple candidate plans from scratch.
The last row of Tab. II summarizes the number and proportion of inferior plans (i.e., out-of-memory or time-limit-exceeded) generated by each method across the three benchmarks. Spark SQL produces a substantial number of inferior plans, particularly on the STACK and JOB benchmarks, primarily due to its limited join optimization capabilities. AutoSteer also produces a number of inferior plans, as its learned optimization strategy tends to favor disabling high-overhead rules. While this can reduce planning time, it often backfires on complex queries. In contrast, AQORA does not generate any inferior plans during test query execution, although a few may arise during training. This highlights the stability and robustness of AQORA’s optimization strategy.
TABLE II: The number of queries showing performance improvement/regression compared with Spark SQL’s default configuration.
JOB ExtJOB STACK
$\delta\in$ Spark Lero AutoSteer AQORA Spark Lero AutoSteer AQORA Spark Lero AutoSteer AQORA
$(0,0.2)$ - 3 1 41 - 4 0 9 - 19 8 22
$(0.2,+\infty)$ - 18 88 20 - 6 0 8 - 51 64 75
(0.2,0)(-0.2,0) - 15 11 41 - 9 8 5 - 19 29 7
$(-\infty,-0.2)$ - 77 13 11 - 5 16 2 - 21 9 6
Failure 10(8.8%) 0 20(17.6%) 0 0 0 0 0 33(30.0%) 0 34(30.1%) 0
VII-B4 Tail Latency
Tail performance is an important indicator of an optimizer’s stability. To provide a comprehensive view, we report the 30th, 60th, 90th, and 99th percentile execution times. Fig. 8 illustrates the end-to-end query latencies of Spark SQL, Lero, AutoSteer, and AQORA at these percentiles after training to convergence. We observe that AQORA consistently outperforms the other methods on all three benchmarks across most percentile levels.
Figure 8: Comparison of execution times at the 30th, 60th, 90th, and 99th percentiles of test queries across three benchmarks for different optimization methods.
On the JOB benchmark, AQORA achieves a 99th percentile latency of 66.86 s, yielding a 99.18% reduction compared to Lero. Notably, the other two baselines fail to produce executable plans at this percentile, underscoring AQORA’s robustness in handling complex queries and its significant advantage in tail latency reduction.
On the Stack benchmark, AQORA consistently outperforms all baselines at the 90th and 99th percentiles. Its raw execution latency at these percentiles is comparable to that of Lero, demonstrating that AQORA can achieve similarly high-quality plans without enumerating multiple candidates.
At lower percentiles, AQORA sometimes underperforms AutoSteer because the latter disables several costly optimization rules. For simpler queries that dominate these percentiles, skipping such rules yields better performance.
VII-B5 Dynamic Evaluation
The first row of Fig. 9 illustrates the performance of AQORA and Lero under data distribution changes. The full IMDb database we use contains data up to 2013. We construct two additional versions, IMDb-1950 and IMDb-1980, by filtering out all movies released after 1950 and 1980, along with their related entries in other tables. IMDb-1950 contains less than 10% of the full IMDb data, while IMDb-1980 retains about 30%[38]. We train the optimizers on these old versions of IMDb and then test them on the complete IMDb database to evaluate their robustness to data evolution. The second row of Fig. 9 shows the results where the optimizers are trained on the training set of JOB and tested on ExtJOB, and vice versa.
Figure 9: The first row shows the performance of AQORA and Lero when trained on IMDb-1950 and IMDb-1980 and tested on the full IMDb dataset. The second row presents the results where the optimizers are trained on the JOB and tested on ExtJOB, and vice versa.
We observe that AQORA consistently outperforms both Spark SQL and Lero under various experimental settings, except for a slight performance drop below Spark SQL when trained on ExtJOB queries using the IMDb-1950 dataset and tested on the full IMDb database. This result demonstrates AQORA’s robustness to dynamic environmental changes. Such robustness mainly stems from two factors: (1) the use of accurate runtime cardinalities enables AQORA to adapt its optimization behavior to the current execution context; and (2) the simplified feature encoding and lightweight network architecture effectively prevent overfitting.
However, both learning-based approaches experience noticeable performance degradation when the execution environment changes. AQORA shows relatively larger drops in raw execution performance. Specifically, Lero’s maximum raw execution degradation across the four settings is 10.80%, 7.86%, 7.53%, and 8.43%, while AQORA’s corresponding drops are 18.04%, 34.93%, 6.99%, and 18.33%. In many cases, AQORA fails to select the optimal plan because the required action was rarely explored during training, whereas Lero can almost always ensure that even an unchosen plan at training at least belongs to its explored set. If AQORA were trained using an offline RL algorithm equipped with a comprehensive experience buffer, this limitation might be mitigated.
VII-C Case Study
In this section, we analyze the results on a per-query basis to find out what AQORA actually achieves in optimizing query plans. In particular, we compare the end-to-end query execution time of learning-based methods with those generated by Spark SQL’s native optimizer. For a query plan $p$ executed by a learning-based method, we calculate its improvement or regression as $\delta=(C_{\text{Spark}}(p)-C_{\text{LQO}}(p))/C_{\text{Spark}}(p)$. Tab. II summarizes the distribution of $\delta$ for each method.
VII-C1 Query Improvement Analysis
As shown in Fig. 10, AQORA achieves up to 89.81%, 69.33%, and 91.83% performance improvement on individual queries from the JOB, ExtJOB, and STACK benchmarks, respectively.
Figure 10: The top 10 queries improved by AQORA in each benchmark compared with Spark SQL’s default configuration.
For example, in JOB query Q16a, AQORA not only enabled the CBO but also refined the plan by prioritizing the join between table title and table movie_keyword, significantly reducing the execution time from over 300 s to just 30.56 s, demonstrating AQORA’s ability to refine CBO-generated plans by improving join orders while balancing optimization overhead and performance gains.
[…]
VII-C2 Query Regression Analysis
While AQORA delivers substantial improvements on many queries, we also observe performance regressions on a small subset. Minor regressions are due to the optimization overhead introduced by AQORA.
Significant regressions stem from limitations of AQORA’s current learning framework, which occasionally leads to suboptimal plan choices (e.g., JOB Q28b). To mitigate this, we introduced several restrictions during training, such as limiting the number of candidate lead tables and capping the maximum number of action steps. These constraints help guide the agent toward more stable decisions and reduce exploration noise. However, such regressions are not entirely avoidable and remain an area for future improvement.
VII-C3 Proportion of bushy plans
Here we analyze the proportion of test queries whose final executed plans adopt a bushy join structure. In our experiments, 22 (19.5%) queries in JOB, 5 (20.1%) in ExtJOB, and 30 (27.3%) in STACK ultimately execute with bushy plans, enabled by the application of CBO actions and leading actions as described in Sec. VI-B1. These results demonstrate that AQORA is capable of effectively exploring bushy join plans to improve query performance.
VII-D Ablation Study
In this ablation study, we will try to replace AQORA’s core components with other possible substitutes to test out the effectiveness of AQORA’s design on ExtJOB. AQORA uses PPO as the default reinforcement learning algorithm, a TreeCNN architecture for both actor and critic and action networks, and an action space that includes CBO actions, leading actions (with only one allowed leaf node), and no-op actions. Learning is conducted through the curriculum process described in Sec.V-B3.
VII-D1 Different Reinforcement Learning Algorithms
To evaluate the impact of different reinforcement learning algorithms on query optimization performance, we compare the PPO algorithm used in AQORA with DQN [39]. As shown in Fig. 11(a), AQORA with PPO surpasses Spark SQL in execution latency significantly earlier during training compared to its DQN-based variant. Also, after training on 2400 queries, the PPO-based model achieves consistently lower execution latency than the DQN-based model. Although DQN benefits from experience replay, it suffers from limitations in environments with large action spaces and non-stationary dynamics.
Figure 11: Query performance on ExtJOB with different substituted modules, under varying numbers of training queries.
VII-D2 Different Neural Network Structures
We tested multiple neural network architectures for the actor and critic networks, with the basic model statistics and optimization overhead summarized in Tab. III. Among them, FCNN is the only model that fails to outperform Spark SQL even after training on 2400 queries. Specifically, LSTM exhibits the lowest optimization overhead (251 ms), making it more efficient during both training and inference. However, TreeCNN demonstrates better convergence behavior and achieves higher overall performance after training, albeit with a slightly increased overhead.
TABLE III: Parameters and optimization overhead of different model structure on JOB and ExtJOB dataset
Model Parameters Optimization overhead
TreeCNN 147506 317 ms per query
LSTM 4800 251 ms per query
QueryFormer 879602 671 ms per query
FCNN 410546 432 ms per query
VII-D3 Different Learning Strategy
To evaluate the effectiveness of our learning strategy design, we conducted an ablation study using two alternative strategies (Fig. 11(c)). In the first strategy, we removed the step limit during optimization while retaining the curriculum learning schedule. In the second strategy, we limited the maximum number of steps but removed the curriculum constraint, allowing AQORA to explore the full action space from the beginning.
Experimental results show that the first strategy performs comparably to the default configuration during the early training phase. However, as the curriculum progresses to its second stage, the model tends to take significantly more steps, making it difficult to repeatedly make correct decisions. The second strategy, despite enforcing a strict step limit, ultimately fails to outperform Spark SQL. This is because, in most cases, only a few targeted corrections are sufficient to reach an efficient plan. Without early-stage guidance, the model explores a large action space too freely, which hinders convergence.
VII-D4 Different Action Space
Although AQORA’s planner extension supports a wide range of action types, we conducted ablation studies to identify which subsets of the action space contribute most to learning efficiency and convergence. Specifically, we experimented with enabling or disabling the broadcast hint, the leading hint, and the CBO toggle. As shown in the results, enabling the broadcast hint often leads the model to generate plans that broadcast excessively large tables during training, resulting in significant performance fluctuations and instability.
When leading actions are disabled, AQORA loses its ability to explicitly manipulate join ordering, thereby reducing its optimization power and leading to suboptimal plans for many complex queries. Finally, disabling the CBO action prevents AQORA from jumping to an entirely different planning path starting from a CBO-generated plan. This restriction increases the difficulty for AQORA to escape local optima and discover globally better execution plans.
VIII Conclusion
We propose AQORA, the first reinforcement learning-based optimizer on Spark SQL that supports execution-time adaptation through four key innovations: (1) realistic feature encoding, (2) stage-level feedback with real-time intervention, (3) automatic strategy adaptation, and (4) seamless integration via Spark’s extension interface. By dynamically constructing bushy joins through node swapping and broadcast tuning, AQORA reduces Spark SQL’s execution latency by up to 70%, while achieving up to 99.8% faster optimization compared to enumeration-based learning query optimizers.